OpenStudy (anonymous):

(03.07) Graph with a line going through points zero comma two point five and four comma two point five. Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph. y = -1 y = 3 y = x -1 y = 3x - 1 @ankitshaw

OpenStudy (anonymous):

(03.07) Graph with a line going through points negative 2 comma zero and 0 comma negative one. Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1

OpenStudy (anonymous):

@ankitshaw

OpenStudy (ankitshaw):

 this was for first question

OpenStudy (anonymous):

i cant see half of what you said

OpenStudy (anonymous):

Graph with a line going through points negative 2 comma zero and 0 comma negative one. Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1 @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

What was your answer for this one?

OpenStudy (anonymous):

y = - 2x + 2

OpenStudy (yanasidlinskiy):

And she said it was wrong?

OpenStudy (anonymous):

yes but i thought it was right

OpenStudy (yanasidlinskiy):

I'm confused here...For what question is this, because I see you have three posted up. Do you mind starting from the first one and working our way out till the last one??

OpenStudy (anonymous):

03.07) Graph with a line going through points zero comma two point five and four comma two point five. Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph. y = -1 y = 3 y = x -1 y = 3x - 1 This was the first one, the answer was -1

OpenStudy (anonymous):

for that i chose C

OpenStudy (yanasidlinskiy):

First of all, do you know what the correct answer is?

OpenStudy (anonymous):

A

OpenStudy (ankitshaw):

A is the correct answer i already solved it above but it wasnt clear to see

OpenStudy (yanasidlinskiy):

Ok..What do you not understand? What is your doubt?

OpenStudy (ankitshaw):

@jadee27p option A that is y = -1 is correct

OpenStudy (anonymous):

she never explained why i got it wrong but that was the first one i got wrong, the second one was this.

OpenStudy (anonymous):

(03.07) Graph with a line going through points negative 2 comma zero and 0 comma negative one. Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1 i chose y = - 2x + 2 my teacher said that was also wrong.

OpenStudy (yanasidlinskiy):

@jadee27p .. Clear this up for me Please. What is actually your question? I don't really undersrand

OpenStudy (anonymous):

i want to know how i got the 2nd one wrong.

OpenStudy (yanasidlinskiy):

So you want to understand why you got this wrong?? Correct? Graph with a line going through points negative 2 comma zero and 0 comma negative one. Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1

OpenStudy (anonymous):

yes

OpenStudy (ankitshaw):

what was the answer you got???

OpenStudy (anonymous):

y = - 2x + 2

OpenStudy (anonymous):

That was my answer and i dont know why i got a 0 for that problem.

OpenStudy (yanasidlinskiy):

Points:(-2,0)(0,-1) The slope of this line is change in y over change in x

OpenStudy (anonymous):

so my answer was correct.

OpenStudy (ankitshaw):

the answer is y= 2x +4

OpenStudy (yanasidlinskiy):

Lost connection again sorry about that @jadee27p

OpenStudy (yanasidlinskiy):

But what was you teachers answer again?

OpenStudy (ankitshaw):

@YanaSidlinskiy the correct answer is y = 2x + 4 damn sure

OpenStudy (yanasidlinskiy):

@ankitshaw Why are you using profanity?

OpenStudy (ankitshaw):

come on i wasnt being rude , but sorry if u felt so

OpenStudy (yanasidlinskiy):

@jadee27p ...Just go with @ankitshaw...Since he thinks it's right!:) Not going to refer to this anymore!:) @ankitshaw I know you weren't being rude, but this website is for helping not using for profanity!

OpenStudy (ankitshaw):

sorry my word choice to express my self was not appropriate ... i apologize

OpenStudy (yanasidlinskiy):

It's ok @ankitshaw ! I forgive you!:)

OpenStudy (ankitshaw):

@YanaSidlinskiy thanx :D

OpenStudy (yanasidlinskiy):

No problem:)